• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 16:00:17 on December 17,1999

    Entry number 31318

    keyword=Shift Summary

    08:00---I_beam =45 uA, Q2=15.0
    Still taking data ( 3He, P_e=0.743Gev/c, P_h=0.0984Gev/c,
    Angle_e=57.911, Angle_h=25.206)

    Start Run 1393

    The T2 rate is about 300 and the T1 is only about 50. Checked the t1 and t2 scintillator hit histograms using the HANA. It seems the t2 events are just backgrounds. Called Bob to look into this problem. He checked those histograms
    and also thought those events are just backgrounds and he said that happened in gammap P experiment too. Maybe for this kinematics the backgrounds are too high.

    08:30---End Run 1393 ( Nevts=0.33M, T1=81K)
    Start Run 1394
    08:50---End Run 1394(Nevts=0.41M, T1=100K)
    MCC stopped the beam to fix the injector problem.

    09:40---Replayed all of the data for run 1394. There are about 1000 elastic events.
    About 1% of the T1 events is elastic. The elastic event rate is only 0.8Hz at
    the beam 45uA.

    Still waiting for the Beam

    10:30---MCC starts sending pulsed beam to Hall A.

    11:50---MCC is still tuning the beam.

    12:00---Start doing shift checklist.
    The GUI for HV monitoring doesn't work. No high voltage display on the
    GUI. Seems it couldn't get those high voltage values.

    Checked the detector histogram. They all look fine.
    14:40---MCC told us the injector people still couldn't find a soultion to send the 70
    uA current. They will just send Hall A about 50 uA beam current until the end of this

    14:50---MCC called they might be able to send us 70uA CW beam.

    15:00---CW Beam comes to Hall A. (67uA)
    Ask MCC to turn on the Fast Feedback . MCC is still trying to keep a stable
    high cauuren.
    Start Run 1395. Raster size is OK.
    15:20---End Run 1395. ( OTR is out)
    Replayed Run 1395, everthing is fine.

    MCC told us they are going to put the OTR in.

    The Beam is 70 uA now. And it looks pretty stable.

    15:21--- Start Run 1396
    Ask MCC to adjust the beam position to (0,0).
    MCC told us the 70uA is beam limit to the target. So they will send us 68uA maximum beam.
    Beam 66uA.
    15:57---Continue data taking